OpenStudy (anonymous):

How do I write and simplify an expression for the area of the square: √8 - √5?

OpenStudy (anonymous):

A = lw = (√8 - √5)(√8 - √5) = √64 - √40 -√40 + √25 = 8 - √40 - √40 + 5 = 13 - √80 and I am now stuck.

OpenStudy (anonymous):